Your company estimates it will have to replace one of its delivery cars five years from now. The replacement is estimated to cost not less than Gh 160,000, which it plans to finance by depositing an equal amount at the end of each month into an account with a return of 0.90% per month. What is the amount to be deposited each month in order to realize the required amount?
1
Expert's answer
2020-06-18T20:22:26-0400

A=P((1+r/m)mt-1)/(r/m)

P=(A*(r/m))/((1+r/m)mt-1)

P=(160000*(0.009/12))/((1+0.009/12)60-1)=120/0.046=2608.7